# Different Victims, Same Layout: Email Visual Similarity Detection for Enhanced Email Protection

### v1

@@ Abstract.
In the pursuit of an effective spam detection system, the focus has often been on identifying known spam patterns either through rule-based detection systems or machine learning (ML) solutions. However, both systems are susceptible to evasion techniques and zero-day attacks that can be achieved at low cost. Therefore, an email that bypassed the defense system once can do it again in the following days, even though rules are updated or the ML models are retrained. The recurrence of failures to detect emails that exhibit layout similarities to previously undetected spam is concerning for customers and can erode their trust in a company. Our observations show that threat actors reuse email kits extensively and can bypass detection with little effort, for example, by making changes to the content of emails. In this work, we propose an email visual similarity detection approach, named Pisco, to improve the detection capabilities of an email threat defense system. We apply our proof of concept to some real-world samples received from different sources. Our results show that email kits are being reused extensively and visually similar emails are sent to our customers at various time intervals. Therefore, this method could be very helpful in situations where detection features that rely on contextual information and keywords are bypassed, an occurrence our observations show happens frequently.
@@ 1. Introduction
Our observations show that email kits are reused extensively in the wild (see Figure 1 as an example), and several emails with layout similarities are observed in our telemetry at different time intervals. We believe this can happen for two main reasons: 1) An old threat actor changing their infrastructure but reusing the same email kit as a few days ago, or 2) Different threat actors using the same email kit. Email kits provide different features that enable threat actors to add additional elements (e.g., logos) to emails. They also allow them to re-paraphrase specific parts of emails to evade detection systems that rely on keywords. Therefore, email kit reuse introduces significant challenges to conventional defense mechanisms such as rule-based detection engines and even ML-based solutions that are not robust against zero-day attacks.
@@ 2.1. Image Capturing
Our approach parses each email in the first step to retrieve its contents in accordance with the RFC822 standard (Force, 2021). Then, email banners are removed, and the email is re-written. Banners are added to emails by difference services, and they serve different purposes. For example, security warnings are typically added in a banner to the message by email security gateways. Once the banner is removed, the message is rendered in the Thunderbird email client via an X virtual framebuffer, known as Xvfb. When fully loaded, the message screenshot is taken and stored for image processing.
@@ 2.2.1. Normalization
In image processing, normalization typically involves modifying pixel values to adhere to a particular range or distribution, often with the goal of enhancing contrast or standardizing intensity values. Common normalization techniques include histogram equalization and contrast stretching, both of which seek to reorganize pixel intensities across a wider spectrum, resulting in a visually appealing image. We rely on the NORM_MINMAX method of the OpenCV library to normalize the emails’ screenshots (on Arrays via OpenCV, 2024).

### v3

@@ Abstract.
In the pursuit of an effective spam detection system, the focus has often been on identifying known spam patterns either through rule-based detection systems or machine learning (ML) solutions that rely on keywords. However, both systems are susceptible to evasion techniques and zero-day attacks that can be achieved at low cost. Therefore, an email that bypassed the defense system once can do it again in the following days, even though rules are updated or the ML models are retrained. The recurrence of failures to detect emails that exhibit layout similarities to previously undetected spam is concerning for customers and can erode their trust in a company. Our observations show that threat actors reuse email kits extensively and can bypass detection with little effort, for example, by making changes to the content of emails. In this work, we propose an email visual similarity detection approach, named Pisco, to improve the detection capabilities of an email threat defense system. We apply our proof of concept to some real-world samples received from different sources. Our results show that email kits are being reused extensively and visually similar emails are sent to our customers at various time intervals. Therefore, this method could be very helpful in situations where detection engines that rely on textual features and keywords are bypassed, an occurrence our observations show happens frequently.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAdoi: 10.1145/3658644.3691381isbn: 979-8-4007-0636-3/24/10
@@ 1. Introduction
Our observations show that email kits are reused extensively in the wild (see Figure 1 as an example), and several emails with layout similarities are observed in our telemetry at different time intervals. We believe this can happen for two main reasons: 1) An old threat actor changing their infrastructure but reusing the same email kit as a few days ago, or 2) Different threat actors using the same email kit. Email kits provide different features that enable threat actors to add additional elements (e.g., logos) to emails. They also allow them to re-paraphrase specific parts of emails to evade detection systems that rely on keywords. Therefore, email kit reuse introduces significant challenges to conventional defense mechanisms such as rule-based detection engines and even ML-based solutions that rely on textual features and are not robust against zero-day attacks.
@@ 2.1. Image Capturing
Our approach parses each email in the first step to retrieve its contents in accordance with the RFC822 standard (Force, 2021). Then, email banners are removed, and the email is re-written. Banners are added to emails by different services, and they serve different purposes. For example, security warnings are typically added in a banner to the message by email security gateways. Once the banner is removed, the message is rendered in the Thunderbird email client via an X virtual framebuffer, known as Xvfb. When fully loaded, the message screenshot is taken and stored for image processing.
@@ 2.2.1. Normalization
In image processing, normalization typically involves modifying pixel values to adhere to a particular range or distribution, often with the goal of enhancing contrast or standardizing intensity values. Common normalization techniques, such as histogram equalization and contrast stretching, reorganize pixel intensities across a wider spectrum, resulting in a visually appealing image. We rely on the NORM_MINMAX method of the OpenCV library to normalize the emails’ screenshots (on Arrays via OpenCV, 2024).